OpenStudy (anonymous):

suppose S is between R and T. use the segment addition posulate to solve for the variable.. so the question is RS= 7y-4 can some one simplify it.. cuase i dont understand

OpenStudy (anonymous):

segment addition postulate is AB+BC=AC

OpenStudy (anonymous):

so it RS=7y-4 ST= y+5 RT= 28 sorry i left thoose out

OpenStudy (anonymous):